Two Fields have the same perimeter one is a square if side 72 m & another is a rectangle of 80 m which plot is greater .
Answer by macston(5194)

perimeter of square=4s=4(72)=288
perimeter of rectangle=2L+2W=same as square, L=80
288=2(80)+2W
288=160+2W subtract 160 from each side
288-160=160+2W-160
128=2W divide each side by 2
128/2=2W/2
64=W Width of rectangle is 64 m
Area of rectangle=L x W=80m x 64m=5120 sq m
Area of square=s x s=72m x 72m=5184 sq m
The square plot is 64 sq m larger than the rectangular plot.